OpenStudy (ria23):

Which expression gives the x-coordinate of the solution of the system? {2x + 3y = 5 {3x + 5y = 7 (Again, it's one big bracket.)

OpenStudy (ria23):

Which expression gives the y-coordinate of the solution of the system? {2x + y = 2 {3x + 4y = -22 and Which expression shows the solution of the system? {5x + 4y = 2 {-2y = 8 + 4x If the chosen answers are needed I'll post those too.

OpenStudy (ria23):

@shiraz14

OpenStudy (shiraz14):

Think I'd need to see the system you are referring to? You might be missing a diagram here - please upload it. Thanks.

OpenStudy (ria23):

I believe it's referring to the system of equations. It didn't give me anything other than this.

OpenStudy (shiraz14):

@Ria23 : I'd need to be away for ~1 hour ... you can direct someone else if you wish. Thanks.

OpenStudy (ria23):

Ok, Thank yhu for yhur help.

OpenStudy (ria23):

@johnweldon1993 Do yhu understand how to wor these out? ;c I'm stuck and I have no idea what to do.

OpenStudy (ria23):

work*

OpenStudy (johnweldon1993):

Well just to see what type of answers they want to choose from...post the choices from the first question...the original topic one above :)

OpenStudy (ria23):

Ok. c:

OpenStudy (johnweldon1993):

Hmm....didn't want to open for me :/

OpenStudy (johnweldon1993):

Well we can at least go through the process of solving for 'x' in the first one... Lets use the elimination method here... 2x + 3y = 5 3x + 5y = 7 We want to eliminate the 'y' ....we want to subtract these equations and make y = 0....so lets multiply the whole top equation by 5...and the whole bottom equation by 3... 5(2x + 3y = 5) ----> 10x + 15y = 25 3(3x + 5y = 7) ----> 9x + 15y = 21 so now our 2 equations are 10x + 15y = 25 9x + 15y = 21 Lets subtract them.... 10x + 15y = 25 - (9x + 15y = 21) ----------------- x = 4

OpenStudy (ria23):

Then plugin x. right?

OpenStudy (johnweldon1993):

Well if you wanted to solve for 'y' after that..yes...but the question above just says expression for the 'x' coordinate of the solution

OpenStudy (ria23):

Dx Oh God I have no idea what I'm doing. The choices are like the other two questions I had before.

OpenStudy (shiraz14):

@Ria23 & @johnweldon1993 : This question does not seem to make any sense to me. It seems that there is a system which needs to be evaluated and the expressions above define its x- & y-coordinates. @Ria23 , please take a screenshot of the actual question and upload it into this thread. Thanks.

OpenStudy (shiraz14):

@Ria23 : Thanks for uploading the screenshot. Could you please check if there was a diagram of a system shown before this question? Some questions relate to diagrams and cannot be addressed unless the actual scenario/context of the question is known.

OpenStudy (ria23):

No, There's no diagram.

OpenStudy (shiraz14):

@Ria23 : Darn, I can't believe it's so simple - the answer to your first question is D.

OpenStudy (ria23):

:o Could yhu explain how yhu came to that answer?

OpenStudy (shiraz14):

Equate the quotient of the matrix dets (as provided in the options) to the deduced value of x in the original equations.

OpenStudy (shiraz14):

@Ria23 : Could you please upload the screenshots of the options for the other 2 questions as well, so we can look at them?

OpenStudy (shiraz14):

I'm logging off now. Goodbye.

OpenStudy (ria23):

Yes.

OpenStudy (ria23):